I need some help on this problem:
I need to solve the system by substitution
x + y = 12
y = 2x
Jamie

Answer by funmath(2933)

Hi Jamie,
Solve the system by substitution
L1) x + y = 12
L2) y = 2x
Substitute 2x into L1 and solve for x:
x+(2x)=12
3x=12
3x/3=12/3
x=4
Substitute 4 into L2 and solve for y:
y=2(4)
y=8
The solution is (x,y)=(4,8)
Happy Calculating!!!